# Cluster deep loci and mirror symmetry

### 26 Feb 2024 v1

@@ 1.5. Back to mirror symmetry
We see that, for this sort of application to mirror symmetry, it is important to understand when the deep locus $\mathcal{D}(\mathcal{A})$ is empty. A necessary condition, and according to Conjecture 1.1 a sufficient one, is that the stabilizer locus, $\mathcal{S}(\mathcal{A})$, is empty. Thus, we will remark throughout out the paper on combinatorial conditions which imply that various stabilizer loci are empty. For example, for $\mathcal{A}(k,n)$, we will show (Corollary 4.44) that the stabilizer locus is empty if and only if $\gcd(k,n)=1$. Combined with Theorem 1.2, this implies that the deep locus is empty for $\mathcal{A}(3,n)$ if $3$ does not divide $n$.
@@ 1.6. Beyond existence: geometry of the deep locus
Albeit pathological from the point of view of algebraic geometry, we are surprised to witness that in all cases each irreducible component of the deep locus $\mathcal{D}(\mathcal{A})$ is again a cluster variety. While this is not the case for an arbitrary cluster variety $\mathcal{A}$, we show in Theorem 6.4 that every irreducible component of the stabilizer locus $\mathcal{S}(X(\beta))$ for an arbitrary double Bott-Samelson variety is again a double Bott-Samelson variety, and so in particular is again a cluster variety. Further, all the intersections of the irreducible components are again double Bott-Samelson varieties and so in particular cluster varieties Thus, the same is true for the deep locus $\mathcal{D}(X(\beta))$ as long as Conjecture 1.1 holds for $X(\beta)$. Similarly, upon the validity of Conjecture 1.1, every irreducible component and every intersection of irreducible components of the deep locus of a positroid variety [43] is a cluster variety. We do not know what is the largest possible generality where one can observe this phenomenon.
For an arbitrary braid variety $X(\beta)$, we show that all irreducible components and their intersections of the analogue of the stabilizer locus for a certain natural action of a torus which in general does not surject onto $\operatorname{Aut}(\beta)$ are again braid varieties. While this result does not immediately fit into the cluster algebra context, we find it interesting that the statement can be interpreted in terms of unlinking the components of the closure of the braid $\beta\Delta^{-1}$, where $\Delta$ is the half-twist braid.
@@ 4.5. Torus actions
In Section 4.7, we will introduce the class of double Bott-Samelson varieties; this is a large class of braid varieties for which the map $T\to\operatorname{Aut}(\mathbb{C}[X(\beta)]$ is surjective. In Section 4.9, we will mention that this also holds for positroid varieties.
@@ 6.2. The deep locus of $X(a,b)$
6.3. The $T$-stabilizer locus of a braid variety and link splitting
Similarly to the proof of Theorem 6.4, we can completely describe the components of the $T$-stabilizer locus of any braid variety $X(\beta)$. We remark, however, that in general the map $T\to\operatorname{Aut}(X(\beta))$ is not surjective, so this does not a priori give a description of the stabilizer locus of $X(\beta)$, see Example 4.29. We will assume that the Demazure product $\delta(\beta)=w_{0}$ is the longest element in $S_{n}$; by [6, Lemma 3.4] we do not lose any generality in doing this. Geometrically, the assumption that $\delta(\beta)=w_{0}$ means that any two strands cross in the braid diagram for $\beta$. The following result is clear.
@@ 6.3. The $T$-stabilizer locus of a braid variety and link splitting
Let $\beta\in\mathrm{Br}_{n}^{+}$ be a positive braid with $\delta(\beta)=w_{0}$. Let $w=\pi(\beta)\in S_{n}$ be the Coxeter projection. Decompose $\{1,\dots,n\}=I_{1}\sqcup I_{2}$ in such a way that both $I_{1}$ and $I_{2}$ are unions of $ww_{0}$-orbits, with $|I_{1}|=a,|I_{2}|=b$. Let $\beta_{1}\in\mathrm{Br}_{a}^{+},\beta_{2}\in\mathrm{Br}_{b}^{+}$ be the braids obtained by considering only the strands indexed by $I_{1}$ and $I_{2}$, respectively. Then, the set of elements in $X(\beta)$ which are stabilized by $T_{I_{1},I_{2}}$ is isomorphic to $X(\beta_{1})\times X(\beta_{2})$.
@@
Note that, since $\delta(\beta)=w_{0}$, we have that any two strands in the braid diagram for $\beta$ cross at least once; it follows that the Demazure products of $\beta_{1},\beta_{2}$ are the longest element in $S_{a}$ and $S_{b}$, respectively. Now the proof follows step by step that of Theorem 6.4, using Lemma 6.12 instead of Lemma 6.2 to conclude. ∎
@@
Since by assumption $z\in X(\beta)$, $w_{0}B_{\beta}(z)$ is upper triangular, so we obtain $t_{w_{0}(i)}=t_{w(i)}$ for every $i$, or equivalently $t_{i}=t_{w_{0}w^{-1}(i)}$ for every $i$. Taking inverses, the result follows. ∎
@@
Let $\beta\in\mathrm{Br}^{+}_{n}$ be a positive braid, with $\delta(\beta)=w_{0}$. Then, the irreducible components of the $T$-stabilizer locus in $X(\beta)$ correspond to the partitions $\{1,\dots,n\}=I_{1}\sqcup I_{2}$, where $I_{1}$ and $I_{2}$ are subsets closed under the action of $ww_{0}$. The component corresponding to $I_{1},I_{2}$ is isomorphic to the braid variety $X(\beta_{1})\times X(\beta_{2})$, where $\beta_{1},\beta_{2}$ are as described in Theorem 6.13.
Here, $X(\beta_{1})\times X(\beta_{2})$ is indeed isomorphic to a braid variety $X(\beta^{\prime})$ for $\beta^{\prime}$ on $n$ strands obtained by permuting the indices of strands in $\beta_{1}\beta_{2}$ by a permutation which sends $I_{1}$ to $\left\{1,\ldots,a\right\}$ and $I_{2}$ to $\left\{a+1,\ldots,a+b\right\}$. By multiplying $\beta^{\prime}$ by a suitable reduced word as in [6, Lemma 3.4], we can find $\beta^{\prime\prime}$ with $\delta(\beta^{\prime\prime})=w_{0}$ and
@@
By construction, the inclusion of a subvariety isomorphic to $X(\beta^{\prime\prime})\cong X(\beta_{1})\times X(\beta_{2})$ into $X(\beta)$ is $T$-equivariant. This induces a map in $T$-equivariant cohomology
@@
By iterating arguments in the proof of Theorem 6.13 and Corollary 6.15, we obtain the following.
Corollary 6.17.
Any set partition $I_{1}\sqcup I_{2}\sqcup\ldots\sqcup I_{k}$ of the set of orbits of $ww_{0}$ (equivalently, of the set of components of the closure of $\beta\Delta^{\pm 1}$) gives rise to a subgroup $T_{I_{1},I_{2},\ldots,I_{k}}\subset(\mathbb{C}^{\times})^{n}$. Let $\beta_{j}$ be the braid obtained by considering only the strands indexed by $I_{j}$, for $j=1,\ldots,k$. Then, the set of elements in $X(\beta)$ which are stabilized by $T_{I_{1},I_{2},\ldots,I_{k}}$ is isomorphic to $X(\beta_{1})\times X(\beta_{2})\times\cdots\times X(\beta_{k})$. In particular, the intersection of sets of points stabilized by two subtori of (×$(\mathbb{C}^{\times})^{n}$ is isomorphic to the product of braid varieties associated with the coarsest common refinement of the corresponding partitions of orbits of $ww_{0}$.
The poset of irreducible components of the $T$-stabilizer locus of $X(\beta)$ and their intersections is thus given by the refinement partial order of partitions of the set of orbits of $ww_{0}$, equivalently, of the set of components of the closure of $\beta\Delta^{\pm 1}$.
Remark 6.18.
Although, as remarked above, the map $T\to\operatorname{Aut}(X(\beta))$ is not in general surjective, it is surjective in the case of positroid varieties, cf. Section 4.9 above. So Corollaries 6.15 and 6.17 can be used to describe the irreducible components and their intersections of the stabilizer locus of positroid varieties, which in particular are again cluster varieties.
@@ 7. Applications to mirror symmetry
This section explains how the previous results on deep loci of cluster varieties fit in the context of Homological Mirror Symmetry (HMS). We use them to formulate a conjectural criterion for when the Fukaya category of $G/P$ is generated by tori (Conjecture 7.1), and verify it for an infinite class of Grassmannians (Theorem 7.2); in such cases this also verifies that HMS holds when choosing as mirror Landau-Ginzburg model the one proposed by Rietsch for $G/P$ [57] (see also [48] for the special case of Grassmannians). For the sake readability, we treat the Fukaya category as a blackbox, and give references where the technical definitions and main foundational results used in the proof are established.

### 05 Mar 2024 v2

@@ 1.5. Back to mirror symmetry
We see that, for this sort of application to mirror symmetry, it is important to understand when the deep locus $\mathcal{D}(\mathcal{A})$ is empty. A necessary condition, and according to Conjecture 1.1 a sufficient one, is that the stabilizer locus, $\mathcal{S}(\mathcal{A})$, is empty. Thus, we will remark throughout the paper on combinatorial conditions which imply that various stabilizer loci are empty. For example, for $\mathcal{A}(k,n)$, we will show (Corollary 4.44) that the stabilizer locus is empty if and only if $\gcd(k,n)=1$. Combined with Theorem 1.2, this implies that the deep locus is empty for $\mathcal{A}(3,n)$ if $3$ does not divide $n$.
@@ 1.6. Beyond existence: geometry of the deep locus
Albeit pathological from the point of view of algebraic geometry, we are surprised to witness that in all cases each irreducible component of the deep locus $\mathcal{D}(\mathcal{A})$ is again a cluster variety. While this is not the case for an arbitrary cluster variety $\mathcal{A}$, we show in Theorem 6.4 that every irreducible component of the stabilizer locus $\mathcal{S}(X(\beta))$ for an arbitrary double Bott-Samelson variety is again a double Bott-Samelson variety, and so in particular is again a cluster variety. Further, all the intersections of the irreducible components are again double Bott-Samelson varieties and so in particular cluster varieties. Thus, the same is true for the deep locus $\mathcal{D}(X(\beta))$ as long as Conjecture 1.1 holds for $X(\beta)$. Similarly, upon the validity of Conjecture 1.1, every irreducible component and every nonempty intersection of irreducible components of the deep locus of a positroid variety [43] is a cluster variety. We do not know what is the largest possible generality where one can observe this phenomenon.
For an arbitrary braid variety $X(\beta)$, we show that all irreducible components of the analogue of the stabilizer locus for a certain natural torus action (which in general does not coincide with that of $\operatorname{Aut}(X(\beta))$) are again braid varieties and their intersections, when not empty, are also braid varieties. While this result does not immediately fit into the cluster algebra context, we find it interesting that the statement can be interpreted in terms of unlinking the components of the closure of the braid $\beta\Delta^{-1}$, where $\Delta$ is the half-twist braid.
@@ 2.1. Cluster algebras
Remark 2.5.
When $Q$ is the empty quiver, or equivalently when $\tilde{B}$ is the empty matrix, the cluster algebra $A$ is declared to simply be $\mathbb{C}$.
@@ 4.5. Torus actions
In Section 4.7, we will introduce the class of double Bott-Samelson varieties; this is a large class of braid varieties for which the map $T\to\operatorname{Aut}(\mathbb{C}[X(\beta)])$ is surjective. In Section 4.9, we will mention that this also holds for positroid varieties.
@@ 6.1. The stabilizer locus for double Bott-Samelson varieties
Iterating the arguments in the proof of Theorem 6.4 and Corollary 6.6 we obtain the following result.
Corollary 6.7.
Let $\beta\in\mathrm{Br}_{n}^{+}$ be a positive braid. Any set partition $I_{1}\sqcup I_{2}\sqcup\dots\sqcup I_{k}$ of $\{1,\dots,n\}$ where each $I_{j}$ is a union of orbits of $\pi(\beta)$ (or equivalently, a set partition $C=C_{1}\sqcup\dots\sqcup C_{k}$ on the set of connected components of the braid closure $\overline{\beta}$ gives rise to a subgroup $T_{I_{1},\dots,I_{k}}\subseteq(\mathbb{C}^{\times})^{n}$. For $j=1,\dots,k$ let $\beta_{j}$ be the braid obtained by considering only the strands indexed by $I_{j}$. Then, the set of elements of $\mathrm{BS}(\beta)$ which are stabilized by $T_{I_{1},\dots,I_{k}}$ is isomorphic to $\mathrm{BS}(\beta_{1})\times\cdots\times\mathrm{BS}(\beta_{k})$. In particular, the intersection of the sets of points stabilized by two subtori of (×$(\mathbb{C}^{\times})^{n}$ is isomorphic to the product of double Bott Samelson varieties associated with the coarsest common refinement of the corresponding partitions of the set $C$.
Note that there may be two different set partitions of $C$ giving rise to the same locus of $\mathrm{BS}(\beta)$ in Corollary 6.7. For example, if $\beta=\sigma_{1}\sigma_{2}\sigma_{2}\sigma_{1}\in\mathrm{Br}_{3}^{+}$ we have that $\overline{\beta}$ consists of three components, $C=\{C_{1},C_{2},C_{3}\}$, so that $C_{2}$ and $C_{3}$ do not intersect. If we set $I_{1}=\{C_{1}\},I_{2}=\{C_{2},C_{3}\}$ and $I^{\prime}_{1}=\{C_{1}\}$, $I^{\prime}_{2}=\{C_{2}\}$, $I^{\prime}_{3}=\{C_{3}\}$, we can see that these two set partitions give rise to the same locus in $\mathrm{BS}(\beta)$.
@@ 6.2. The deep locus of $X(a,b)$
6.3. The $T$-stabilizer locus of a braid variety
Similarly to the proof of Theorem 6.4, we can completely describe the components of the $T$-stabilizer locus of any braid variety $X(\beta)$. We remark, however, that in general the map $T\to\operatorname{Aut}(X(\beta))$ is not surjective, so this does not a priori give a description of the stabilizer locus of $X(\beta)$, see Example 4.29. We will assume that the Demazure product $\delta(\beta)=w_{0}$ is the longest element in $S_{n}$; by [6, Lemma 3.4] we do not lose any generality in doing this. The following result is clear.
@@ 6.3. The $T$-stabilizer locus of a braid variety
Let $\beta\in\mathrm{Br}_{n}^{+}$ be a positive braid with $\delta(\beta)=w_{0}$. Let $w=\pi(\beta)\in S_{n}$ be the Coxeter projection. Decompose $\{1,\dots,n\}=I_{1}\sqcup I_{2}$ in such a way that both $I_{1}$ and $I_{2}$ are unions of $ww_{0}$-orbits, with $|I_{1}|=a,|I_{2}|=b$. Let $\beta_{1}\in\mathrm{Br}_{a}^{+},\beta_{2}\in\mathrm{Br}_{b}^{+}$ be the braids obtained by considering only the strands indexed by $I_{1}$ and $I_{2}$, respectively. Then:
  1. (1)
    If $\delta(\beta_{1})\in S_{I_{1}}$ and $\delta(\beta_{2})\in S_{I_{2}}$ are the longest elements in their respective symmetric groups, then the set of element of $X(\beta)$ which are stabilized by $T_{I_{1},I_{2}}$ is isomorphic to $X(\beta_{1})\times X(\beta_{2})$.
  2. (2)
    Else, the set of elements in $X(\beta)$ which are stabilized by $T_{I_{1},I_{2}}$ is empty.
@@
[…]
Remark 6.16.
Note that $\mathrm{BS}(\beta)\cong X(\Delta\beta)$, cf. Lemma 4.34 and any two strands cross in a braid diagram for $\Delta\beta$. Thus, Case (2) of Theorem 6.15 cannot happen for double Bott-Samelson varieties and hence the statement of Theorem 6.4 is easier than that of Theorem 6.15.
Remark 6.17.
In general, Case (2) of Theorem 6.15 can happen. For example, take $\beta=\sigma_{1}\sigma_{1}\sigma_{2}\sigma_{2}\sigma_{1}\sigma_{1}$, cf. Example 4.29. We have $w=\pi(\beta)=e$ and $ww_{0}=w_{0}=[3,2,1]$. Thus, the only non-trivial decomposition of $\{1,2,3\}$ into unions of $ww_{0}$-orbits is $I_{1}=\{1,3\}$, $I_{2}=\{2\}$. But the first and third strands do not cross in a braid diagram for $\beta$, so $\delta(\beta_{1})=e$ and we are in Case (2) of Theorem 6.15. Note that this is consistent with Example 4.29, which implies that the $T$-stabilizer locus of $X(\beta)$ is empty.
Note that we obtain the following corollary of Theorem 6.15, which generalizes Corollary 4.38.
Corollary 6.18.
Let $\beta\in\mathrm{Br}^{+}_{n}$ be a braid such that any two strands cross in its braid diagram. Then $T$ acts freely on $X(\beta)$ if and only if the braid closure $\overline{\beta\Delta}$ is a knot, that is, $\pi(\beta)w_{0}$ is an $n$-cycle.
@@
Since by assumption $z\in X(\beta)$, $w_{0}B_{\beta}(z)$ is upper triangular, we obtain that $t_{w_{0}(i)}=t_{w(i)}$ for every $i$, or equivalently $t_{i}=t_{w_{0}w^{-1}(i)}$ for every $i$. Taking inverses, the result follows. ∎
@@
Let $\beta\in\mathrm{Br}^{+}_{n}$ be a positive braid, with $\delta(\beta)=w_{0}$. Then, the irreducible components of the $T$-stabilizer locus in $X(\beta)$ are given by partitions $\{1,\dots,n\}=I_{1}\sqcup I_{2}$, where $I_{1}$ and $I_{2}$ are subsets closed under the action of $\pi(\beta)w_{0}$ for which (1) of Theorem 6.15 occurs. The component corresponding to $I_{1},I_{2}$ is isomorphic to the braid variety $X(\beta_{1})\times X(\beta_{2})$, where $\beta_{1},\beta_{2}$ are as described in Theorem 6.15.
Note that two distinct partitions $\{1,\dots,n\}$ as in Corollary 6.20 can give the same irreducible component of the $T$-stabilizer locus. For example, if $\beta=\Delta$ then $X(\Delta)$ is a point (which is its own stabilizer locus), $\pi(\Delta)w_{0}=e$, and any partition of $\{1,\dots,n\}$ gives $X(\Delta)$.
In Corollary 6.20 $X(\beta_{1})\times X(\beta_{2})$ is indeed isomorphic to a braid variety $X(\beta^{\prime})$ for $\beta^{\prime}$ on $n$ strands obtained by permuting the indices of strands in $\beta_{1}\beta_{2}$ by a permutation which sends $I_{1}$ to $\left\{1,\ldots,a\right\}$ and $I_{2}$ to $\left\{a+1,\ldots,a+b\right\}$. By multiplying $\beta^{\prime}$ by a suitable reduced word as in [6, Lemma 3.4], we can find $\beta^{\prime\prime}$ with $\delta(\beta^{\prime\prime})=w_{0}$ and
@@
Similarly to Corollary 6.7, iterating arguments in the proof of Theorem 6.15 and Corollary 6.20, we obtain the following.
Corollary 6.21.
Let $\beta\in\mathrm{Br}^{+}_{n}$ be a positive braid, with $\delta(\beta)=w_{0}$. Any set partition $I_{1}\sqcup I_{2}\sqcup\ldots\sqcup I_{k}$ of $\{1,\dots,n\}$ into sets which are stable under the action of $\pi(\beta)w_{0}$ (equivalently, of the set of components of the closure of $\beta\Delta^{\pm 1}$) gives rise to a subgroup $T_{I_{1},I_{2},\ldots,I_{k}}\subset(\mathbb{C}^{\times})^{n}$. Let $\beta_{j}$ be the braid obtained by considering only the strands indexed by $I_{j}$, for $j=1,\ldots,k$. Then, the set of elements in $X(\beta)$ which are stabilized by $T_{I_{1},I_{2},\ldots,I_{k}}$ is either:
  1. (1)
    Isomorphic to $X(\beta_{1})\times X(\beta_{2})\times\cdots\times X(\beta_{k})$, if, for all $j$, $\delta(\beta_{j})$ is the longest element in $S_{I_{j}}$.
  2. (2)
    Empty, otherwise.
In particular, the intersection of sets of points stabilized by two subtori of (×$(\mathbb{C}^{\times})^{n}$ is either empty or isomorphic to the product of braid varieties associated with the coarsest common refinement of the corresponding partitions of orbits of $ww_{0}$.
@@
If the map $T\to\operatorname{Aut}(X(\beta))$ is surjective, then Corollaries 6.20 and 6.21 describe the components of the stabilizer locus in $X(\beta)$ and their intersections. In particular, these describe the components of the stabilizer locus of positroid varieties and their intersections, which in particular are again cluster varieties when nonempty.
6.4. Link splitting
Let $\beta\in\mathrm{Br}_{n}^{+}$ and assume that $\{1,\dots,n\}=I_{1}\sqcup I_{2}$ is a decomposition into $\pi(\beta)w_{0}$ stable subsets such that we are in Case (1) of Theorem 6.15. By construction, the inclusion of a subvariety isomorphic to $X(\beta^{\prime\prime})\cong X(\beta_{1})\times X(\beta_{2})$ into $X(\beta)$ is $T$-equivariant. This induces a map in $T$-equivariant cohomology
@@ 7. Applications to mirror symmetry
This section explains how the previous results on deep loci of cluster varieties fit in the context of Homological Mirror Symmetry (HMS). We use them to formulate a conjectural criterion for when the Fukaya category of $G/P$ is generated by tori (Conjecture 7.1), and verify it for an infinite class of Grassmannians (Theorem 7.2); in such cases this also verifies that HMS holds when choosing as mirror Landau-Ginzburg model the one proposed by Rietsch for $G/P$ [57] (see also [48] for the special case of Grassmannians). For the sake of readability, we treat the Fukaya category as a blackbox, and give references where the technical definitions and main foundational results used in the proof are established.